OpenStudy (anonymous):

Evaluate the limit by first recognizing the sum as the Riemann sum for a function defined on the interval [0,1]: lim as n approaches infinity (1/n)((sqrt(1/n)+(sqrt(2/n)+(sqrt(3/n)+...............+(sqrt(n/n))

OpenStudy (anonymous):

i have no idea how to start this question

OpenStudy (zarkon):

\[\int\limits_{0}^{1}\sqrt{x}dx\]

OpenStudy (anonymous):

would i have to use reimann sums to get the answer right? Or just to to find the function

OpenStudy (anonymous):

(1/n)((sqrt(1/n)+(sqrt(2/n)+(sqrt(3/n)+...............+(sqrt(n/n)) \[(1/n)*\sum_{1}^{n}\sqrt[2]{(i/n)}=0.6667\]

OpenStudy (anonymous):

\[(1/n)[\sqrt{(1/n)}+\sqrt{(2/n)}+\sqrt{(3/n)}+...+\sqrt{(n/n)}=0.66667\]

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\sqrt{x} dx=0.66666667=2/3\]